OpenStudy (anonymous):

Please help, will give medal! David kicks a soccer ball off the ground and in the air, with an initial velocity of 35 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 17.9 feet 18.2 feet 18.7 feet 19.1 feet

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

What do you need? answer or logic?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

for right now you can just give me the answer

OpenStudy (loser66):

if the ball is kicked from the ground then s =0, right? with v = 35, your equation is H(t) = -16t^2 +35t hence the maximum value is the vertex of the graph we find t-coordinate by \(t=\dfrac{-b}{2a}=\dfrac{-35}{2*(-16)}=1.0937\) replace that t into the function to find \( H(t) =-16 (1.0937^2+35*(1.0937)= 19.1\)